# Intrinsic training dynamics of deep neural networks

### v1

@@ 4.2 Infinitely deep linear networks
We next consider the limit when $L\rightarrow+\infty$ of deep linear residual networks with parameters $U_{k}=\mathrm{Id}_{n}+\mathcal{A}_{\frac{k}{L}}$, and thus focus on the analysis of the parameter $\theta=(\mathcal{A}_{s})_{s\in[0,1]}$, where $\mathcal{A}_{s}\in\mathbb{R}^{n\times n}$: this corresponds to studying linear neural ODEs (introduced by Chen et al. (2018)). Remarkably, our theoretical approach continues to apply in this regime, and yields a closed-form formula for the metric. In this section, we thus study the dynamics of parameters $\theta(t)\in\mathcal{X}$ where $\mathcal{X}$ corresponds to the Banach space $(\mathcal{C}^{1}([0,1],\mathbb{R}^{n\times n}),\|\cdot\|_{\mathcal{C}^{1}})$ where $\|f\|_{\mathcal{C}^{1}}:=\max\!\bigl{\{}\|f\|_{\infty},\,\|f^{\prime}\|_{\infty}\bigr{\}}$, and such that the trajectory $t\mapsto\theta(t)=(\mathcal{A}_{s}(t))_{s\in[0,1]}$ is the solution of the gradient flow on $\ell(\theta)$, given by the (family of coupled) ODE
@@
$\forall s\in[0,1],\quad\frac{\partial\mathcal{A}_{s}}{\partial t}(t)=-\,\mathfrak{g}_{s}(t),\quad\text{with}\quad\mathfrak{g}_{s}(t):=\frac{\partial\ell}{\partial\mathcal{A}_{s}}\Bigl{(}\theta(t)\Bigr{)}\in\mathbb{R}^{n\times n},$ (16)
@@
where we assume that the loss function $\ell:\mathcal{X}\mapsto\mathbb{R}$ is such that $\theta\mapsto(\frac{\partial\ell}{\partial\mathcal{A}_{s}}(\theta))_{s\in[0,1]}$ is locally Lipschitz on $\mathcal{X}$ (to ensure by the Cauchy-Lipschitz theorem that indeed there exists a unique maximal solution $\theta(\cdot)\in\mathcal{C}^{1}([0,T),\mathcal{X})$ of (16) with a given $\theta(0)$).
@@
Now as an infinite-depth analog of the product $Z_{L}=U_{L}\ldots U_{1}$, given any $\theta\in\mathcal{X}$ we consider $s\in[0,1]\mapsto Z_{s}=Z_{s}[\theta]\in\mathbb{R}^{n\times n}$ the unique global solution (as $\theta=(\mathcal{A}_{s})_{s\in[0,1]}\in\mathcal{X}$) of the state equation
$\frac{\mathrm{d}}{\mathrm{d}s}Z_{s}=\mathcal{A}_{s}Z_{s},\quad Z_{0}=\mathrm{Id}_{n}.$ (17)
@@
The analog to ˜2.1 is to assume that $\ell(\theta)=f\bigl{(}Z_{s=1}\bigr{)}$ with $f\in\mathcal{C}^{1}$, and we now want to know if it is possible to rewrite the dynamic $\frac{\partial Z_{s=1}}{\partial t}(t)$ as an intrinsic dynamic that only depends on $Z_{s=1}(t)$ and the initialization $\theta(0)$.
The following proposition (see Appendix˜L for a proof) gives a set of conserved functions during all trajectories of (16).
@@
For any $\theta\coloneqq(\mathcal{A}_{s})_{s\in[0,1]}\in\mathcal{X}$, we define for any $s\in[0,1]$ the function $\mathbf{h}_{s}:\theta\mapsto\mathcal{A}_{s}^{\prime}+\mathcal{A}_{s}^{\prime\top}+[\mathcal{A}_{s}^{\top},\mathcal{A}_{s}]\in\mathbb{R}^{n\times n},$ where we denote $\mathcal{A}_{s}^{\prime}\coloneqq\frac{\mathrm{d}}{\mathrm{d}s}\mathcal{A}_{s}$. Then for any $s\in[0,1]$, one has for any $t$: $\mathbf{h}_{s}(\theta(t))=\mathbf{h}_{s}(\theta(0))$, where $\theta(t)$ is the maximal solution of (16) with initialization $\theta(0)$.
Moreover, the following theorem (see Appendix˜M for a proof) shows that for relaxed balanced initializations, the evolution of $Z_{1}(t)=Z_{s=1}(t)$ is entirely described by $Z$ and the initialization.
@@
with $\gamma(s)\coloneqq(1-s)\psi_{1}(1)-\psi_{1}(1-s)-s\psi_{2}(1)+\psi_{2}(s)$, where $\psi_{1}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u$ and $\psi_{2}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u$. In particular, if $\lambda(\cdot)\equiv 0$ (balanced-condition), then $\gamma(\cdot)\equiv 0$.
In a sense, this theorem captures the infinite-depth limit ($L\rightarrow+\infty$) of Theorem˜4.6, while offering the key advantage of an explicit closed-form expression for the associated metric.
@@ Conclusion
Aknowledgments
The work of G. Peyré was supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir" program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute) and by the European Research Council (ERC project WOLF). The work of R. Gribonval was supported by the AllegroAssai ANR-19-CHIA-0009 project of the French Agence Nationale de la Recherche and by the SHARP ANR project ANR-23-PEIA-0008 in the context of the France 2030 program.
Appendix A More comments on Example˜2.5
In example (2.5), when $\lambda\rightarrow+\infty$ $K_{\theta_{0}}/\lambda\to I_{m}$ (the Euclidean metric). When $\lambda=0$, one has $K_{\theta_{0}}(z)=\|z\|I_{m}+\frac{zz^{\top}}{\|z\|}$ for every $z\neq 0$, and in particular, by the uniqueness result in the Cauchy-Lipschitz theorem, 0 is reachable only if $z_{0}=0$. When $\lambda\rightarrow$-$\infty$, $K_{\theta_{0}}/|\lambda|\to\frac{zz^{\top}}{\|z\|^{2}}$. See https://github.com/sibyllema/Intrinsic_training_dynamic for numerical illustrations of these different behaviors.
Appendix B Proof of Theorem˜2.14
@@
Appendix C Proof of Theorem˜3.3
See 3.3
@@
We first show $(i)\implies(ii)$. We assume $(i)$. Observe that given any $\theta\in\mathbb{R}^{D}$(8) is equivalent (by rank theorem) to
@@
By smoothness of $\phi$ and $\mathbf{h}$, if (18) holds at $\theta_{0}$ is also holds in a whole neighborhood $U$ of $\theta_{0}$. By the implicit function theorem, denoting $F(\theta)\coloneqq(\phi(\theta),\mathbf{h}(\theta))$, it implies that $\theta$ can be expressed on $UU$ as $\theta=F^{-1}(\phi(\theta),\mathbf{h}(\theta))=\Gamma(\phi(\theta),\mathbf{h}(\theta))$.
@@
Appendix D Proof of Proposition˜3.7.
See 3.7
@@
$(i)\implies(ii)$ is direct consequence of the proof of (Marcotte et al., 2023, Proposition 3.7).
We now show $(ii)\implies(iii)$. Let us assume $(ii)$. We fix $\theta_{0}$. Then by assumption on $U\ni\theta_{0}$, $\theta=\Gamma(\mathbf{h}(\theta),\phi(\theta))$, and by using Proposition˜3.7 one has $\ker\partial\phi(\theta)\cap\ker\partial\mathbf{h}(\theta)=\{0\}$ on $\mathcal{M}_{\theta_{0}}\cap U$. Thus (18) holds on a open neighborhood $O$ of $\theta_{0}$. As $\mathbf{h}$ are conservation laws for $\phi$ on $U$, one has on $O\cap U$ that $\mathbb{W}_{\phi}(\theta)=D-\text{rank}\partial h(\theta)$.
But as $\langle\nabla h(\theta),\nabla\phi_{i}(\theta)\rangle=\langle\nabla h(\theta),\nabla\phi_{j}(\theta)\rangle=0\implies\langle\nabla h(\theta),[\nabla\phi_{i},\nabla\phi_{j}](\theta)\rangle=0$, one has necessarly $\mathrm{dim}\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})\leq D-\text{rank}\partial h(\theta_{0})$ and as one also has $D-\text{rank}\partial h(\theta_{0})=\mathrm{dim}\mathbb{W}_{\phi}(\theta_{0})\leq\mathrm{dim}\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})$ then one has $\mathbb{W}_{\phi}(\theta_{0})=\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})$. This holds for any $\theta_{0}$, which concludes the proof. ∎
Appendix E Proof of Theorem˜3.8.
See 3.8
@@
Appendix F Proof of Corollary˜3.9
We first prove the following proposition (recall that $\mathbb{W}_{\phi}$ is defined in (9) and below).
@@
If $\phi:\mathbb{R}^{D}\to\mathbb{R}^{d}$ is polynomial, then there exists a dense open set $\Theta$ of $\mathbb{R}^{D}$ such that for any $\theta\in\Theta$, $\mathrm{dim}\mathbb{W}_{\phi}(\theta)=\max_{\theta^{\prime}\in\mathbb{R}^{D}}\mathrm{dim}\mathbb{W}_{\phi}(\theta^{\prime})$.
@@
Denote $M\coloneqq\max_{\theta^{\prime}\in\mathbb{R}^{D}}\mathrm{dim}\mathbb{W}_{\phi}(\theta^{\prime})\in\mathbb{N}$. Considering $\theta_{0}\in\mathbb{R}^{D}$ such that $\mathrm{dim}\mathbb{W}_{\phi}(\theta_{0})=M$, there exists distinct indices $i_{1},\cdots,i_{M}$ such that the vectors $\nabla\phi_{i_{j}}(\theta_{0})\in\mathbb{R}^{D},1\leq j\leq M$ are linearly independent. There also exists a set $I$ of $M$ coordinates such that the restriction of these vectors to $I$ remains linearly independent. The function $\theta\mapsto\eta(\theta)\coloneqq\text{det}[(\nabla\phi_{i_{1}}(\theta))_{I},\cdots,(\nabla\phi_{i_{M}}(\theta)_{I}]$ is a polynomial on $\mathbb{R}^{D}$ with $\eta(\theta_{0})\neq 0$, hence the set $\mathcal{Z}$ of its zeros is a closed negligeable set of $\mathbb{R}^{D}$. Thus the open dense subset of $\mathbb{R}^{D}$ defined by $\Theta\coloneqq\mathbb{R}^{D}\setminus\mathcal{Z}$ satisfies: for all $\theta\in\Theta$: $\mathrm{dim}\mathbb{W}_{\phi}(\theta)=M$. ∎
See 3.9

### v2

@@ 3.1.2 Infinitely deep linear networks
We next consider the limit when $L\rightarrow+\infty$ of deep linear residual networks with parameters $U_{k}=\mathrm{Id}_{n}+\mathcal{A}_{\frac{k}{L}}$, and thus focus on the analysis of the parameter $\theta=(\mathcal{A}_{s})_{s\in[0,1]}$, where $\mathcal{A}_{s}\in\mathbb{R}^{n\times n}$, which corresponds to a linear neural ODEs model (introduced by Chen et al. (2018)). Remarkably, our theoretical approach still applies in this regime, and yields a closed-form formula for the metric. We thus study the dynamics of parameters $\theta(t)\in\mathcal{X}$ where $\mathcal{X}$ corresponds to the Banach space $(\mathcal{C}^{1}([0,1],\mathbb{R}^{n\times n}),\|\cdot\|_{\mathcal{C}^{1}})$ where $\|f\|_{\mathcal{C}^{1}}:=\max\!\bigl\{\|f\|_{\infty},\,\|f^{\prime}\|_{\infty}\bigr\}$, and such that the trajectory $t\mapsto\theta(t)=(\mathcal{A}_{s}(t))_{s\in[0,1]}$ is the solution of the gradient flow on $\ell(\theta)$, given by the (family of coupled) ODEs
@@
$\forall s\in[0,1],\quad\tfrac{\partial\mathcal{A}_{s}}{\partial t}(t)=-\,\mathfrak{g}_{s}(t),\quad\text{with}\quad\mathfrak{g}_{s}(t):=\tfrac{\partial\ell}{\partial\mathcal{A}_{s}}\Bigl(\theta(t)\Bigr)\in\mathbb{R}^{n\times n},$ (15)
@@
where we assume that the loss function $\ell:\mathcal{X}\mapsto\mathbb{R}$ is such that $\theta\mapsto(\frac{\partial\ell}{\partial\mathcal{A}_{s}}(\theta))_{s\in[0,1]}$ is locally Lipschitz on $\mathcal{X}$ (to ensure by the Cauchy-Lipschitz theorem that indeed there exists a unique maximal solution $\theta(\cdot)\in\mathcal{C}^{1}([0,T),\mathcal{X})$ of equation 15 with a given $\theta(0)$).
@@
As an infinite-depth analog of $Z_{L}=U_{L}\ldots U_{1}$, given any $\theta\in\mathcal{X}$ we consider $s\in[0,1]\mapsto Z_{s}=Z_{s}[\theta]\in\mathbb{R}^{n\times n}$ the unique global solution (as $\theta=(\mathcal{A}_{s})_{s\in[0,1]}\in\mathcal{X}$) of the state equation
$\tfrac{\mathrm{d}}{\mathrm{d}s}Z_{s}=\mathcal{A}_{s}Z_{s},\quad Z_{0}=\mathrm{Id}_{n}.$ (16)
@@
The analog to ˜2.1 is to assume that $\ell(\theta)=f\bigl(Z_{s=1}\bigr)$ with $f\in\mathcal{C}^{1}$, and we now want to know if it is possible to rewrite the dynamics $\frac{\partial Z_{s=1}}{\partial t}(t)$ as an intrinsic dynamic that only depends on $Z_{s=1}(t)$ and the initialization $\theta(0)$. The following proposition (see Appendix˜L for a proof) gives a set of conserved functions during all trajectories of equation 15.
@@
For any $s\in[0,1]$, consider $\mathbf{h}_{s}:\theta\coloneqq(\mathcal{A}_{s})_{s\in[0,1]}\in\mathcal{X}\mapsto\mathcal{A}_{s}^{\prime}+\mathcal{A}_{s}^{\prime\top}+[\mathcal{A}_{s}^{\top},\mathcal{A}_{s}]\in\mathbb{R}^{n\times n},$ where we denote $\mathcal{A}_{s}^{\prime}\coloneqq\frac{\mathrm{d}}{\mathrm{d}s}\mathcal{A}_{s}$. Then for any $s\in[0,1]$, one has for any $t$: $\mathbf{h}_{s}(\theta(t))=\mathbf{h}_{s}(\theta(0))$, where $\theta(t)$ is the maximal solution of equation 15 with initialization $\theta(0)$.
Moreover, the following theorem (see Appendix˜M for a proof) shows that for relaxed balanced initializations, the evolution of $Z_{1}(t)=Z_{s=1}(t)$ is intrinsic as it depends only on $Z_{1}$ and the initialization $\theta(0)$:
@@
with $\gamma(s)\coloneqq(1-s)\psi_{1}(1)-\psi_{1}(1-s)-s\psi_{2}(1)+\psi_{2}(s)$, where $\psi_{1}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u$ and $\psi_{2}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u$. If $\lambda(\cdot)\equiv 0$ (balanced-condition), then $\gamma(\cdot)\equiv 0$.
In a sense, this theorem captures the infinite-depth limit ($L\rightarrow+\infty$) of Theorem˜3.11, while offering the key advantage of an explicit closed-form expression for the associated metric.
@@ Conclusion
Acknowledgments
The work of G. Peyré was supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute) and by the European Research Council (ERC project WOLF). The work of R. Gribonval was supported by the AllegroAssai ANR-19-CHIA-0009 project of the French Agence Nationale de la Recherche and by the SHARP ANR project ANR-23-PEIA-0008 in the context of the France 2030 program. The authors thank Stephan Thomassé, Gilles Blanchard and Laurent Jacques for discussions on rank properties of DAGs that helped simplify the proof of Proposition˜F.7 in Section˜F.2.
Appendix A A table that summarizes which parametrizations can be used to analyze which type of neural network, along with the corresponding results.
Network type Parametrization $\bm{\phi_{\text{Lin}}}$ $\bm{\phi_{\text{ReLU}}}$
Linear network IMP: only for relaxed balanced $\theta_{0}$ Dimension: $d^{\prime}\bm{\ll}D$ IMP: for all $\theta_{0}$ $d^{\prime}=D-k\bm{\approx}D$
DAG ReLU N/A IMP: for all $\theta_{0}$ $d^{\prime}=D-k\bm{\approx}D$
Appendix B More comments on Example˜2.5
In example equation 2.5, when $\lambda\rightarrow+\infty$ $K_{\theta_{0}}/\lambda\to\mathrm{Id}_{m}$ (the Euclidean metric). When $\lambda=0$, one has $K_{\theta_{0}}(z)=\|z\|\mathrm{Id}_{m}+\frac{zz^{\top}}{\|z\|}$ for every $z\neq 0$, and in particular, by the uniqueness result in the Cauchy-Lipschitz theorem, 0 is reachable only if $z_{0}=0$. When $\lambda\rightarrow$-$\infty$, $K_{\theta_{0}}/|\lambda|\to\frac{zz^{\top}}{\|z\|^{2}}$. See https://github.com/sibyllema/Intrinsic_training_dynamic for numerical illustrations of these different behaviors.
Appendix C Proof of Theorem˜2.14
@@ Acknowledgments
Appendix D Proof of Theorem˜2.17
See 2.17
@@
We first show $(i)\implies(ii)$. We assume $(i)$. Observe that given any $\theta\in\mathbb{R}^{D}$equation 8 is equivalent (by rank theorem) to
@@
By smoothness of $\phi$ and $\mathbf{h}$, if equation 17 holds at $\theta_{0}$ it also holds in a whole neighborhood $U$ of $\theta_{0}$. By the implicit function theorem, denoting $F(\theta)\coloneqq(\phi(\theta),\mathbf{h}(\theta))$, it implies that $\theta$ can be expressed on $U$ as $\theta=F^{-1}(\phi(\theta),\mathbf{h}(\theta))=\Gamma(\phi(\theta),\mathbf{h}(\theta))$.
@@
Appendix E Proof of Proposition˜2.21
See 2.21
@@
$(i)\implies(ii)$ is a direct consequence of the proof of (Marcotte et al., 2023, Proposition 3.7).
We now show $(ii)\implies(iii)$. Let us assume $(ii)$. We fix $\theta_{0}$. Then by assumption on $U\ni\theta_{0}$, $\theta=\Gamma(\mathbf{h}(\theta),\phi(\theta))$, and by using Proposition˜2.21 one has $\ker\partial\phi(\theta)\cap\ker\partial\mathbf{h}(\theta)=\{0\}$ on $\mathcal{M}_{\theta_{0}}\cap U$. Thus equation 17 holds on an open neighborhood $O$ of $\theta_{0}$. As $\mathbf{h}$ are conservation laws for $\phi$ on $U$, one has on $O\cap U$ that $\mathrm{dim}(\mathbb{W}_{\phi}(\theta))=D-\text{rank}\partial\mathbf{h}(\theta)$.
But as for any conservation law $h$ for $\phi$, one has $\langle\nabla h(\theta),\nabla\phi_{i}(\theta)\rangle=\langle\nabla h(\theta),\nabla\phi_{j}(\theta)\rangle=0\implies\langle\nabla h(\theta),[\nabla\phi_{i},\nabla\phi_{j}](\theta)\rangle=0$, one has necessarly $\mathrm{dim}\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})\leq D-\text{rank}\partial\mathbf{h}(\theta_{0})$ and as one also has $D-\text{rank}\partial\mathbf{h}(\theta_{0})=\mathrm{dim}\mathbb{W}_{\phi}(\theta_{0})\leq\mathrm{dim}\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})$ then one has $\mathbb{W}_{\phi}(\theta_{0})=\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})$. This holds for any $\theta_{0}$, which concludes the proof. ∎
Appendix F Proof of Theorem˜3.1, Theorem˜3.2, Corollary˜3.3 and Corollary˜3.4
F.1 Proof of Theorem˜3.1
See 3.1
@@
F.2 Proof of Theorem˜3.2.
To prove Theorem˜3.2, we introduce the notion of prunable DAGs and characterize them, before showing (see Corollary˜F.11) that the rank of $\partial\phi(\theta)$ is almost everywhere equal to that of $\partial\phi(\mathbf{1}_{D})$
F.2.1 Prunable DAGs
Definition F.1.
Given a DAG $G=(V,E)$ with vertices $V$ and edges $E$ and, for each edge $i\in E$ we denote $G^{\prime}_{i}=(V,E^{\prime}_{i})$ with $E^{\prime}_{i}:=E\setminus\{i\}$ the smaller DAG obtained by removing edge $i$. The input neurons of $G$ (neurons without any antecedent) are always also input neurons of $G^{\prime}_{i}$, and similarly for the output neurons of $G$ (neurons without any successor), and we say that $G$ is prunable at edge $i$ if $G^{\prime}_{i}$ has no additional input/output neurons. The graph $G$ is said to be prunable everywhere (resp. somewhere/nowhere) if it is prunable at every (resp. at some/at no) edge $i$.
Recalling that, by definition, a path in a graph is a sequence of consecutive edges starting from an input neuron and ending at an output neuron, we denote $\mathcal{P}_{i}\coloneqq\{p:p\ni i\}$ the set of all paths of $G$ containing an edge $i\in E$, and $\mathcal{P}_{i}^{c}$ the other paths of $G$.
Lemma F.2.
The set $\mathcal{P}_{i}^{c}$ is always a subset of the paths of $G^{\prime}_{i}$. Moreover, if $G$ is prunable at edge $i$, then $\mathcal{P}_{i}^{c}$ is exactly the set of all paths of $G^{\prime}_{i}$.
Proof.
If $p\in\mathcal{P}_{i}^{c}$ then its edges are distinct from $i$, hence $p$ is a path of $G^{\prime}_{i}$. Conversely, when $G$ is prunable at $i$, $G^{\prime}_{i}$ has the same input and output neurons as $G$, hence if $p$ is a path of $G^{\prime}_{i}$ it is a path of $G$ with edges in $E^{\prime}_{i}$, i.e. with no edge equal to $i$, i.e. $p\in\mathcal{P}_{i}^{c}$. ∎
The following straightforward characterizations will be useful.
Lemma F.3.
A graph $G$ is prunable at edge $i$ if, and only if, the output neuron of edge $i$ has at least another incoming edge, and the input neuron of edge $i$ has at least another outgoing edge.
Corollary F.4.
A graph $G$ is prunable everywhere if, and only if, every non-output neuron has at least two successors, and every non-input neurons has at least two antecedents.
Corollary F.5.
A graph $G$ associated to a multilayer perceptron (MLP) of depth $L$ with layer widths $N_{\ell}$, $0\leq\ell\leq L$, is prunable everywhere if, and only if, $N_{\ell}\geq 2$ for every layer $\ell$ (including input and output layers).
F.2.2 Rank properties
Lemma F.6.
[…]
Proof.
By definition of the component $\phi_{p}$ of $\phi=(\phi_{p})_{p}$ for each path $p$ as a product Gonon et al. (2024), one has:
$\frac{\partial\phi_{p}}{\partial{\theta}_{i}}(\theta){\theta}_{i}=\phi_{p}(\theta)1_{\{i\in p\}}.$
In particular for $\theta=\mathbf{1}_{D}$ one gets:
$S\coloneqq\partial\phi(\mathbf{1}_{D})=(1_{\{i\in p\}})_{p,i},$
and thus:
$D_{\phi(\theta)}S=\partial\phi(\theta)D_{\theta},$ (18)
where $D_{\phi(\theta)}=\mathrm{diag}(\phi(\theta))$ and $D_{\theta}=\mathrm{diag}(\theta)$. Note that equation 18 holds for each $\theta\in\mathbb{R}^{D}$.
Finally, as all coordinates of $\theta$ are in $\{$$\mathbb{R}\backslash\{0\}$, both $D_{\phi(\theta)}$ and $D_{\theta}$ are invertible and thus $\mathrm{rank}(S)=\mathrm{rank}(\partial\phi(\theta))$, which concludes the proof. ∎
@@ F.2.2 Rank properties
For any DAG ReLU architecture, one has $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))\geq D-m$ with $m$ the number of hidden neurons and $D$ the number of parameters.
@@
Without loss of generality we prove the result on a connected DAG. Indeed, if the DAG has $C$ connected components, $G_{k}$, $1\leq k\leq C$, each with $D_{k}$ edges and $m_{k}$ hidden neurons, then as $\sum_{k}D_{k}=D$ and $\sum_{k}m_{k}=m$ the result for the whole graph follows from the result on each component: up to row/column permutations, it is easy to check that the matrix $\partial\phi(\mathbf{1}_{D})$ is the block-diagonal concatenation of the corresponding matrices for each component, hence its rank is $\sum_{k}(D_{k}-m_{k})=D-m$.
Assume that $G$ is a connected DAG, and denote $S^{\prime}\coloneqq\partial\phi(\mathbf{1}_{D})[F,:]$, the incidence matrix of $G$ restricted to $F$, the set of all full paths (i.e., paths starting from an input neuron and ending at an output neuron). As a direct consequence of the definition of $S^{\prime}$, one gets: $\mathrm{range}(\partial\phi(\mathbf{1}_{D}))\geq\mathrm{range}(S^{\prime})$. Thus, to conclude the proof it is enough to show that $\mathrm{range}(S^{\prime})=D-m$.
We now show that $\mathrm{range}(S^{\prime})=D-m$. We will use the known fact (see (Gleiss et al., 2003, Proposition 6)) that in a strongly connected222a directed graph where for every nodes $\mu$$\nu$ there exists a directed path starting at $\mu$ and ending at $\nu$ directed graph $G^{\prime}=(E^{\prime},V^{\prime})$, the rank of the cycle matrix $C$ is $|E^{\prime}|-|V^{\prime}|+1$, where the cycle matrix has as many rows as the number of cycles in $G^{\prime}$ and each row has binary entries indicating which edge (indexing the columns of $C$) belongs to the corresponding cycle. To exploit this fact, we complete $G$ (as a DAG, it has no cycle!) into a strongly connected graph $G^{\prime}$ with cycle matrix $C$ satisfying $\mathrm{rank}(S^{\prime})=\mathrm{rank}(C)$, and $|E^{\prime}|=|E|+|V_{\mathrm{out}}|+|V_{\mathrm{in}}|+1$, $|V^{\prime}|=|V|+2$, so that $\mathrm{rank}(S^{\prime})=\mathrm{rank}(C)=|E|-|V|+|V_{\mathrm{out}}|+|V_{\mathrm{in}}|=D-m$ as claimed, since $m=|V|-(|V_{\mathrm{out}}|+|V_{\mathrm{in}}|)$. Above, we denoted $V_{\mathrm{in}}$ (resp. $V_{\mathrm{out}}$) the set of input (resp. output) neurons of $G$.
The construction is as follows: we add two fresh neurons $\mu,\nu$ ($V^{\prime}:=V\cup\{\mu,\nu\}$); the first one is a "source" neuron with an outgoing edge towards each input neuron in $V_{\mathrm{in}}$; the second is a "sink" neuron with an incoming edge from each output neuron in $V_{\mathrm{out}}$; another edge is added $\nu\to\mu$. It is not difficult to check that cycles of $G^{\prime}$ are in bijection with paths in $G$: each cycle $c$ of $G^{\prime}$ passes through exactly one input neuron and one output neuron, and its restriction to the edges of $G$ is precisely a path $p$ between these neurons. Vice-versa, every path $p$ in $G$, starting at $\gamma\in V_{\mathrm{in}}$ and ending at $\gamma^{\prime}\in V_{\mathrm{out}}$ is uniquely prolongated into a cycle of $G^{\prime}$ by adding the three edges $\gamma^{\prime}\to\nu$, $\nu\to\mu$, and $\mu\to\gamma$. With a slight abuse of notation we will thus identify $c$ and $p$.
Equipped with this construction we first show that $S^{\prime}=C[:,E]$, so that $\mathrm{rank}(S^{\prime})\leq\mathrm{rank}(C)$. Indeed, $S^{\prime}$ has binary entries $S^{\prime}[p,e]=\mathbf{1}_{p\ni e}$, so that $S^{\prime}[p,e]=C[c,e]$ for every path and $e\in E$.
To conclude, we only need to show $\mathrm{rank}(C)\leq\mathrm{rank}(S^{\prime})$. For this we prove that each column $C[:,e^{\prime}]$, $e^{\prime}\in E^{\prime}\setminus E$, is in the span of $S^{\prime}$. Denoting $\mathtt{out}(\gamma)$ (resp. $\mathtt{in}(\gamma)$ the set of outgoing (resp. incoming) edges of neuron $\gamma\in V$, and $\mathtt{suc}(\gamma)$ (resp. $\mathtt{ant}(\gamma)$ the set of successors (resp. antecedent) neurons of $\gamma$, i.e. neurons connected via an outgoing (resp. incoming) edge, we distinguish three cases:
  • $e^{\prime}=\mu\to\gamma$ with $\gamma\in V_{\mathrm{in}}$ some input neuron; a cycle $c$ contains this edge $e^{\prime}$ if, and only if, the corresponding path $p$ of $G$ starts at the input neuron $\gamma$, or equivalently if there exists some neuron $\gamma^{\prime}\in\mathtt{suc}(\gamma)$ such that $p$ contains the edge $\gamma\to\gamma^{\prime}$. As $p$ can only pass through at most one of the outgoing edges of $\gamma$, we obtain $C[c,e^{\prime}]=\sum_{e\in\mathtt{out}(\gamma)}S^{\prime}[p,e]$, and since this holds for every $c$ we get $C[:,e^{\prime}]=\sum_{e\in\mathtt{out}(\gamma)}S^{\prime}[:,e]$, showing indeed that $C[:,e^{\prime}]$ belongs to the span of $S^{\prime}$;
  • $e^{\prime}=\gamma\to\nu$, with $\gamma\in V_{\mathrm{out}}$; similarly, $C[:,e^{\prime}]=\sum_{e\in\mathtt{in}(\gamma)}S^{\prime}[:,e]$;
  • $e^{\prime}=\nu\to\mu$: this edge appears in all cycles hence $C[:,e^{\prime}]=\mathbf{1}$. Since we can partition the cycles according to which edge $e^{\prime\prime}=\mu\to\gamma$, $\gamma\in V_{\mathtt{in}}$ they contain, we have $\mathbf{1}=\sum_{\gamma\in V_{\mathtt{in}}}C[:,\mu\to\gamma]$, and since the first case have established that such $C[:,\mu\to\gamma]$ belong to the span of $S^{\prime}$, so does $C[:,e^{\prime}]=\mathbf{1}$.
Corollary F.8.
For any DAG ReLU architecture, one has $\mathrm{rank}(\partial\phi_{\texttt{ReLU}}(\mathbf{1}_{D}))=D-m$ with $m$ the number of hidden neurons and $D$ the number of parameters.
Before showing this result, we first establish that the $m$ "canonical" conservation laws (Le and Jegelka, 2022) associated to each hidden neuron $1\leq j\leq m$ are independent on a set that contains $\mathbf{1}_{D}$.
Lemma F.9.
For any DAG ReLU architecture, the "canonical" conservation laws obtained in Le and Jegelka (2022) associated to each hidden neuron $1\leq j\leq m$, expressed (denoting $\theta_{j}$ the bias of neuron $j$, $\mathtt{succ}(j)/\mathtt{ant}(j)$ its outgoing/incoming edges, and $\theta_{j}$ the corresponding weights) as
$h_{j}(\theta):=\sum_{i\in\mathtt{succ}(j)}\theta_{i}^{2}-\sum_{i\in\mathtt{ant}(j)}\theta_{i}^{2}-\theta_{j}^{2}$ (19)
are independent in the neighborhood of each parameter $\theta$ such that $\nabla h_{j}(\theta)\neq 0$ for all $j$ (so in particular around any $\theta\in\Theta$ with $\Theta$ the set of Theorem˜3.2).